OpenStudy (anonymous):

y^2-49/(y-7)(y-7) / 5y+35/y^2-7y

OpenStudy (anonymous):

supposed to be a division sign not an X!

OpenStudy (austinl):

\[\frac{(y^2-49)}{(y-7)(y-7)} \div \frac{(5y+35)}{(y^2-7y)}\]

OpenStudy (austinl):

Ok! So for rational divition, you are merely multiplying by the reciprocal. Think of it like this, \[\frac{a}{b} \div \frac{c}{d} = \frac{a}{b} \times \frac{d}{c}\]

OpenStudy (campbell_st):

I'd do some factoring as the 1st step and when you divide by a fraction the rule is flip and multiply... or find the reciprocal and multiply so you get \[\frac{(y - 7)(y + 7)}{(y - 7)(y - 7)} \times \frac{y( y - 7)}{5(y + 7)}\] all you need to do is cancel common factors.

OpenStudy (austinl):

So, you would then go about this problem much like the last one that we worked on.

OpenStudy (anonymous):

thank you both!! so then is it just y/5 ?

OpenStudy (austinl):

Yes, it should be.

OpenStudy (anonymous):

you're a saint, thanks!!

OpenStudy (austinl):

No problem, you have any more questions?

OpenStudy (anonymous):

haha well I have 9 questions like this left to answer so i'm sure i will be needing help! i'll post more as I keep going! :)

OpenStudy (austinl):

Ok, if you need help just type @austinL in a response.